OpenStudy (anonymous):

solve for X (x+9)(x+8)(x+7)=0

OpenStudy (anonymous):

(x+9)(x+8)(x+7)=0 x+9=0 x=-9 x+8=0 x=-8 x+7=0 x=-7

OpenStudy (anonymous):

thnx